OpenStudy (anonymous):

find two numbers which their add is 10 and their product is the max.

myininaya (myininaya):

So we have x+y=10 and we want to maximize P=xy

myininaya (myininaya):

Solve x+y=10 for either x or y and then write P in terms of one variable

myininaya (myininaya):

Now can you tell me do you know calculus or do we need to do this algebraically ?

myininaya (myininaya):

Because we can do it both ways. The cal way is easiest though. Less work!

OpenStudy (anonymous):

we are not suppose to use calculus

myininaya (myininaya):

Cool stuff! Ok so can you follow what I have asked you to do so far?

myininaya (myininaya):

You can solve x+y=10 for either x or y by subtracting one of them on both sides

OpenStudy (anonymous):

yeah i've already done it and i think i got it!!

OpenStudy (anonymous):

but thanks anyway!! = )

myininaya (myininaya):

\[x=10-y\] \[P=xy\] \[P=(10-y)y\] \[P=10y-y^2\] So you got this far?

myininaya (myininaya):

Great forking! I'm glad you got it! :)

OpenStudy (anonymous):

yeah n' i got x=5 n' Y=5 and xy= 25

myininaya (myininaya):

Yep yep! :)